DonaldTrumpofQC:

Radical forms:

DonaldTrumpofQC:

\begin{array}{l}\text{Question}\\ \text{Given } x>0 \text{, simplify } \sqrt[4]{x^{32}} \text{ completely.}\end{array}

Spectrum:

I'm actually learning this in Algebra II. Do you have any idea of where to start?

DonaldTrumpofQC:

@spectrum wrote:
I'm actually learning this in Algebra II. Do you have any idea of where to start?
somewhat.

Spectrum:

Okay well since \(x\) doesn't contain a coefficient, we don't have to worry about rooting it. You always divide the exponent attached to \(x\) by the degree of the root. So in order to get your answer, you simply divide the 32 by 4. Which leaves you with \(x ^{8}\)

Spectrum:

Hope this helps! Enjoy your Friday. \(\color{hotpink}{\heartsuit}\)

jhonyy9:

@donaldtrumpofqc wrote:
\[\begin{array}{l}\text{Question}\\ \text{Given } x>0 \text{, simplify } \sqrt[4]{x^{32}} \text{ completely.}\end{array}\]
do you know this ? \[^{\sqrt[3]{x}}= x^{\frac{ 1 }{ 3 }}\]

Jasonisyours:

this was so last year